OpenStudy (anonymous):

T=2π√(l/g). Show that this pendulum equation is dimensionally consistent.

OpenStudy (unklerhaukus):

\[T=2\pi\sqrt\frac lg\]\[T[\text s]=2\pi\sqrt\frac {l[\text m]}{g[\text m/\text s^2]}\]\[[\text s]=\sqrt\frac {[\text m]}{[\text m/\text s^2]}\]